# Landauer Principle and Thermodynamics of Computation

### v1

@@ X Conclusion and Future Direction
The analysis of the cost (like thermodynamic cost, operational cost) while executing a process has been one of the central points of attraction for physicists as well as computer scientists Aifer et al. (2023); Coles (2023). In physics, it is the thermodynamic cost of the process that plays a critical role, whereas in computer science, it is the computational cost. The analysis of the thermodynamic computational cost for a process was initiated from the physical Church-Turing thesis Kolchinsky (2023), where they conveyed that every computational process is physical. Various approaches have been considered for the analysis of the different computational processes. With the advent of modern statistical theory, the research in this area got boosted, yet it still remains a challenging task to explore thermodynamic computational cost (aka Landauer principle) for OQS, many-body physics, and at the quantum phase transition (QPT) Sachdev (1999); Vojta (2003); Osborne and Nielsen (2002); Heyl (2018); Sen et al. (2005); Prabhu et al. (2011) point. QPT takes place in a quantum many-body system Bandyopadhyay et al. (2021); Sur and Ghosh (2020); Fetter and Walecka (2012); Tasaki (2020); De Chiara and Sanpera (2018); Mukherjee et al. (2007); Bose (2003); Dutta (2015); Gómez et al. (1996); Ganahl et al. (2012); Fukuhara et al. (2013); Subrahmanyam (2004); Iyoda and Sagawa (2018); Tian et al. (2013) which happens at absolute zero temperature as a result of quantum fluctuations. The analysis of the minimum cost of heat dissipation of the erasure process in QPT is an open domain to explore.
In recent times, the applications of these principles are not just restricted to the field of physics or computer science but are also prevalent in other fields like chemical networks Chen et al. (2014); Soloveichik et al. (2008); Murphy et al. (2018); Qian and Winfree (2011), molecular biology Prohaska et al. (2010); Benenson (2012), and even in neurology Laughlin (2001); Balasubramanian et al. (2001).
@@
To get a better understanding of the bond between thermodynamics and the computational process, further investigation in this direction is still required. For example, in the case of a finite automaton, one can investigate the maximum thermodynamic cost that is required to accept a language for automata. Also, one can calculate the minimal cost for any deterministic finite automaton. One can also work on developing a theory to analyze the non-deterministic finite automata in terms of thermodynamics. Models to describe the complex Turing machine, and also network theory from the thermodynamic viewpoint, are an open area of research.
@@ XI Acknowledgement
P.C., A.M. would like to thank Nilakanta Meher and Saikat Sur of the Weizmann Institute of Science for their valuable inputs and suggestions.

### v2

@@ X Conclusion and Future Direction
Both the thermodynamic and computational cost of a process have been of prime interests for physicists as well as computer scientists Aifer et al. (2023); Coles (2023). While the physicists are more concerned with the thermodynamic cost a, the computer scientists mostly care about computational cost. The analysis of the thermodynamic computational cost for a process surfaced from the physical Church-Turing thesis Kolchinsky (2023), where it was conveyed that every computational process is physical. Various approaches have been considered for the analysis of thermodynamic cost of the different computational processes. In this article we have reviewed the recent findings in these line in both the quantum and classical regime as well as provided an overview on the foundation and frameworks of thermodynamics of computation. In recent times, the applications of these principles are not just restricted to the field of physics or computer science but are also prevalent in other fields like chemical networks Chen et al. (2014); Soloveichik et al. (2008); Murphy et al. (2018); Qian and Winfree (2011), molecular biology Prohaska et al. (2010); Benenson (2012), and even in neurology Laughlin (2001); Balasubramanian et al. (2001).
@@
To get a better understanding of the bond between thermodynamics and the computational process, further investigations are still required. For example, in the case of a finite automaton, it is tantalizing to investigate the upper bound thermodynamic cost that is required to accept a language for automata. Also, one can calculate the minimal cost for any deterministic finite automaton. Developing a theory to analyze the non-deterministic finite automata with thermodynamic tools could be promising. Models to describe the complex Turing machine, and also network theory from the thermodynamic viewpoint, are open areas of research.
While modern statistical and quantum information theory have advanced our understanding of the thermodynamic cost of computation, applying these insights to many-body physics, especially near quantum phase transitions (QPTs) Sachdev (1999); Vojta (2003); Osborne and Nielsen (2002); Heyl (2018); Sen et al. (2005); Prabhu et al. (2011) remains a significant challenge. QPTs in a quantum many-body system Bandyopadhyay et al. (2021); Sur and Ghosh (2020); Fetter and Walecka (2012); Tasaki (2020); De Chiara and Sanpera (2018); Mukherjee et al. (2007); Bose (2003); Dutta (2015); Gómez et al. (1996); Ganahl et al. (2012); Fukuhara et al. (2013); Subrahmanyam (2004); Iyoda and Sagawa (2018); Tian et al. (2013) occur at zero temperature due to quantum, rather than thermal fluctuations and are characterized by critical phenomena such as diverging correlation lengths and long-range entanglement. These features complicate the application of standard thermodynamic frameworks and raise open questions about how the minimal heat dissipation, governed by LP, behaves near criticality. Some compelling future direction involves relaxing the assumption of thermal reservoirs. These athermal characteristics could significantly modify the minimal energetic cost of erasure in general as well as near the QPT. Another fundamental challenge lies in understanding the behavior of LB as one approaches absolute zero, where the third law of thermodynamics and quantum effects become increasingly relevant. Additionally, pre-existing correlations between the system and a finite-sized reservoir may influence the erasure process in ways not captured by conventional formulations. Though there have been few explorations along these directions as we have reviewed, further explorations are essential for a deeper understanding of information thermodynamics in realistic, quantum-coherent environments.
@@ XI Acknowledgement
P.C., A.M. would like to thank Nilakanta Meher and Saikat Sur of the Weizmann Institute of Science for their valuable inputs and suggestions. A.M. acknowledges partial support from the Anusandhan National Research Foundation, Government of India, under grant number ANRF/ECRG/2024/003836/PMS.